a coffee shop decides to blend a coffee that sells for 12$ per pound with a coffee that sells for $9 per pound to produce a blend that will sell for $10 a pound. How much of each should be used to yield 20 pound o the new blend?
Answer by lwsshak3(11628)

a coffee shop decides to blend a coffee that sells for 12$ per pound with a coffee that sells for $9 per pound to produce a blend that will sell for $10 a pound. How much of each should be used to yield 20 pound o the new blend?
***
let x=amount of $12/pound coffee
20-x=amount of $9/pound coffee
12x+9(20-x)=10*20
12x+180-9x=200
3x=20
x=20/3
amount of $12/pound coffee should be used=20/3 lbs
amount of $9/pound coffee should be used=40/3 lbs